# Lattice simulations of axion-U(1) inflation: gravitational waves, magnetic fields, and black holes

### v1

@@ 4 Gravitational waves
In this section, we discuss the production of GWs during axion-U(1) inflation. The interaction between the axion and the gauge field results in the generation of GWs, as the produced gauge field has a nonzero transverse traceless component in its energy-momentum tensor. Additionally, scalar field spatial fluctuations produced by the gauge field also contribute to the generation of GWs. GWs are defined as the transverse traceless part of the tensor perturbations ($h_{ij}^{\mathrm{TT}}$) of the metric. Including these tensor perturbations, the metric is given by,
@@
The evolution of these tensor perturbations is governed by the Einstein equation. By linearizing the Einstein equation, we obtain;
@@
Here, the ellipsis represents the irrelevant terms proportional to $g_{ij}$. The energy density of the produced GWs is given by,
@@
The density fraction of GWs at the present epoch is given by,
$\Omega_{\mathrm{GW}}h^{2}=\frac{\rho_{\mathrm{GW}}}{\rho_{c}}\Big{|}_{0}=\frac{\rho_{\mathrm{GW}}}{\rho_{r}}\Big{|}_{e}\Omega_{r}h^{2}\,.$ (4.5)
Here, $\Omega_{r}h^{2}$ represents the radiation density fraction at the present epoch and $\rho_{\mathrm{GW}}/\rho_{r}|_{e}$ denotes the ratio of the GW energy density to the total energy density at the time when the GW energy density reaches saturation.
@@
Further in this section, we study the production of GWs in the axion-gauge field scenario, considering both weak and strong coupling cases. The characteristic current frequency of these produced GWs is in GHz range and is determined by the horizon scale at the end of inflation, see equation (5.3). In figure 3, we show the time evolution of the gauge field energy density normalized by the total energy density in the left panel and the density fraction of the GWs in the right panel. The curves correspond to different coupling strengths; $\alpha/f=30$ (dotted violet), 50 (dashed blue), 60 (dot-dashed blue), 75 (solid orange), and 90 (dot-dashed red). The vertical lines, in corresponding colors, represent the end of inflation for each case. The horizontal solid gray line indicates the bound on the GW energy density from the $\Delta N_{\mathrm{eff}}$ constraints from the CMB. The dashed gray curve represents the sensitivity of the future CMB-S4 mission. Our key findings regarding the constraints on the coupling strength between axion and gauge fields from the CMB bound on $\Delta N_{\mathrm{eff}}$ can be summarized as follows.
@@
  • As shown in figure 1, for $\alpha/f=75$ and 90, once the energy density of the gauge field becomes comparable to the kinetic energy of the axion field, the velocity of the axion decreases. This reduction in axion velocity leads to a slower growth rate of the gauge field compared to the pre-backreaction stage. Consequently, the production rate of GWs also slows down, as shown in figure 3. The amplitude of the generated GWs remains below the bound obtained using Planck data for these cases up until the end of inflation.
@@
  • Previous studies [33, 34] indicated that the coupling strength above $70$ leads to an overproduction of GWs, violating the $\Delta N_{\mathrm{eff}}$ bound from the Planck data of CMB. This bound is for the case of chaotic potential, which is also considered for this case. However, we find that this bound is relaxed in the strong coupling case when gauge field backreaction becomes significant during inflation. The specific details of this bound depend on the duration of GW production after the end of inflation. As evident from figure 3, the produced GWs for $\alpha/f=75$ may violate the bound from the Planck data on the GW amplitude, but this occurs 2–3 $e$-folds after the end of inflation. In our analysis, we have not considered the interaction of the gauge field with other standard model particles. For large coupling strengths, this interaction becomes important and may lead to instantaneous reheating, as discussed in section 4.2 of ref. [43]. Considering these effects would lead to the decay of the gauge field into other standard model species, further affecting GW production. If the gauge field energy density decays significantly within 2–3 $e$-folds after inflation, the produced GW amplitude will be smaller and may well be within the bound obtained from the Planck data (shown by the solid horizontal gray line in figure 3).
@@
    Figure 4: In this figure, we show the time evolution of $\rho_{g}/\rho$ and $\Omega_{GW}h^{2}$ for $\alpha/f=75$ for run D (solid yellow curves) and for a run (dotted black curves) analogous to run D in which the contribution of the $\phi$ term in the total energy-momentum tensor, given in equation(4.3), is neglected in the calculation of the GW energy density.
@@ 5 Magnetogenesis
There is indirect evidence of magnetic fields in the intergalactic medium, inferred from the non-detection of secondary GeV photons from blazars [67, 68, 69, 70]. The presence of such intergalactic magnetic fields motivates the study of their generation in the early universe, potentially during inflation [71, 72, 73, 74, 75, 76, 77, 78, 79] or a first-order phase transition [80, 81, 82, 83, 84], for reviews see ref. [85, 86, 87]. In this section, we study the generation of magnetic fields within the axion-U(1) inflation scenario and calculate their present day strength for different coupling strengths, $\alpha/f$.
@@ 5.1 Gauge-field production
First, we study the evolution of the electric and magnetic components of the gauge fields during inflation in our simulations. The root mean square value of the hypermagnetic field and its coherence length are given by,
@@
Similar definitions are used for the electric part of the gauge fields. Furthermore, assuming the universe evolves in a radiation-dominated phase from the end of our simulations, we can calculate the magnetic field strength today using the following expression:
$B_{\mathrm{rms}}\big{|}_{0}$ $=4.2\,{\rm\mu G}\,\sqrt{\frac{{\int d\log{k}\cdot P_{B}}}{\rho}}\,\,\sqrt{\frac{g_{0}}{3.36}}\left(\frac{B_{\rm CMB}}{3.3~{}\rm\mu G}\right)\,.$ (5.2)
Here, $\rho$ represents the total energy density of the universe at the end of our simulations, $g_{0}$ represents the total number of relativistic degrees of freedom of the energy density today, and $B_{\rm CMB}$ is the magnetic field strength corresponding to the CMB energy density.
In figure 5, we demonstrate the evolution of the electric component with solid lines and the magnetic component with dashed lines for coupling strengths $\alpha/f=30$ (green curve), $50$ (red curve), $60$ (orange curve), $75$ (blue curve), and $90$ (yellow curve). In the right panel, we illustrate the evolution of the coherence length for each coupling strength. The coherence length scale is normalized by the length scale corresponding to the Hubble horizon. The length scale corresponding to the Hubble horizon at the end of inflation is given by
@@
$\frac{a_{e}}{a_{0}}=\left(\frac{g_{0s}}{g_{rs}}\right)^{1/3}\frac{T_{0}}{T_{r}}=2.66\times 10^{-29}\left(\frac{g_{0s}}{3.94}\frac{106.75}{g_{rs}}\right)^{1/3}\left(\frac{106.75}{g_{r}}\right)^{1/4}\frac{T_{0}}{2.73\text{K}}\left(\frac{10^{-6}m_{\rm pl}}{H}\right)^{1/2},$ (5.4)
where $g_{rs}$ and $g_{0s}$ denote the effective degrees of freedom in the entropy density at the end of inflation and the present epoch, respectively. We estimate the reheating temperature, $T_{r}$, by assuming instantaneous reheating using $3H^{2}m_{\text{pl}}^{2}=(\pi^{2}/30)g_{r}T_{r}^{4}$.
Figure 5: The evolution of the electric and magnetic components of the gauge field (left panel) and their coherence lengths (right panel) is shown for different values of $\alpha/f$. The coherence lengths are normalized by the length scale corresponding to the Hubble horizon.
@@ 5.2 Evolution of gauge fields post inflation
After the end of inflation, the universe is dominated by the gauge field in the strong coupling regime of axion-U(1) system. Due to the dominance of the gauge field in the total energy budget, the effective equation of state resembles that of radiation dominance. We assume that reheating in this model occurs through the interaction of the gauge field with other Standard Model fields, as described in [93, 43]. Therefore, we can consider the universe as being radiation-dominated just after the end of inflation.
As discussed earlier, the gauge fields have a blue spectrum at the end of inflation and peaks at a scale smaller than the Hubble horizon size; see figure 5. We assume that the gauge fields remain unaffected from the end of inflation until the epoch when the universe’s conductivity becomes much larger than the Hubble parameter. Once this epoch is reached, the electric component of the gauge field shortened out within a Hubble time scale, and the magnetic field becomes frozen into the rest of the plasma. After this epoch, the magnetic field interacts with the plasma via the Lorentz force and transfers some of its energy to the plasma. As a result, the magnetic field strength decreases. This nonlinear processing of the magnetic field begins when the Alfvén time becomes smaller than the Hubble time, which implies that the corresponding scale factor is $a_{\textrm{nl}}\equiv a_{H}(\mathcal{H}~{}L_{c}/V_{A})$. Here, $a_{H}$ corresponds to the scale factor at the horizon entry for a particular scale, and $V_{A}=\sqrt{B^{2}/(2(\rho+p)}$ is the Alfvén velocity.
@@
A $35$ $6.2\times 10^{-17}$ $1.4\times 10^{-3}$ $8.4\times 10^{-14}$ $8.5\times 10^{-16}$ $7.1\times 10^{-3}$ $3.7\times 10^{-14}$
B $50$ $4.6\times 10^{-16}$ $8.7\times 10^{-3}$ $3.4\times 10^{-14}$ $9.0\times 10^{-15}$ $9.9\times 10^{-2}$ $1.0\times 10^{-14}$
C $60$ $1.7\times 10^{-15}$ $3.0\times 10^{-2}$ $1.8\times 10^{-14}$ $1.8\times 10^{-14}$ $9.9\times 10^{-2}$ $1.0\times 10^{-14}$
D $75$ $1.2\times 10^{-14}$ $1.9\times 10^{-1}$ $7.2\times 10^{-15}$ $1.7\times 10^{-14}$ $1.5\times 10^{-1}$ $8.1\times 10^{-15}$
E $90$ $1.4\times 10^{-14}$ $1.7\times 10^{-1}$ $7.6\times 10^{-15}$
Table 2: $B_{\mathrm{rms}}|_{0}$ and $L_{c}^{\mathrm{NL}}|_{0}$ represent the present-day magnetic field strength and its coherence length obtained for each run, respectively. $B_{\rm bound}$ denotes the lower bound on the magnetic field strength based on blazar observations, where we assume $B_{\rm bound}=10^{-17}{\rm G}\left(L/0.1\text{Mpc}\right)^{1/2}$.
In the nonlinear processing regime, the magnetic field strength decays as $(a_{m}/a_{nl})^{-1/3}$, and the coherence length (which corresponds to the peak scale in our case) increases as $(a_{m}/a_{nl})^{2/3}$ for the fully helical case. After incorporating the nonlinear processing of the fields, the magnetic field strength and its coherence length at the present epoch is given by,
$B_{\mathrm{rms}}\big{|}_{0}$ $=2.03\times 10^{-14}\,\,{\rm G}\,\,\sqrt{\frac{{\int d\log{k}\cdot P_{B}}}{\rho}}\,\,\sqrt{\frac{g_{0}}{3.36}}\left(\frac{B_{\rm CMB}}{3.3~{}\rm\mu G}\right)r^{1/3}\,,$ (5.5)
@@
$L_{c}^{\mathrm{NL}}=0.8~{}{\rm pc}~{}(\mathcal{H}L_{c})\left(\frac{10^{-6}m_{\rm pl}}{H}\right)^{1/2}\,r^{-2/3}.$ (5.6)
@@
Here $r=\max(1,\mathcal{H}~{}L_{c}/V_{A})$ and $B_{\rm CMB}$ denotes the magnetic field strength corresponding to the CMB energy density. We show the time evolution of $\mathcal{H}L_{c}/V_{A}$ in appendix E for the runs given in table 1. After using the above expressions, we obtained the magnetic field strength and its coherence length corresponding to each coupling strength and the values are given in table 2. For the values given in columns 2 and 3 of table 2, we use $({\int d\log{k}\cdot P_{B}}/\rho)^{1/2}$ and $\mathcal{H}L_{c}$ at the end of inflation in equation (5.5) and equation (5.6), respectively. However, for the values given in columns 4 and 5, we use the value of $({\int d\log{k}\cdot P_{B}}/\rho)^{1/2}$ and $\mathcal{H}L_{c}$ one $e$-fold after inflation. In table 1, $B_{\rm bound}$ denotes the lower bound of magnetic field strength obtained from the blazars observation. Here we assume $B_{\rm bound}=10^{-17}{\rm G}\,(L_{C}^{\mathrm{NL}}/0.1{\rm Mpc})^{1/2}$. From the values given in the table 2, we conclude that the magnetic field strength obtained in the axion-U(1) inflation model for coupling strengths $\alpha/f\geq 60$ can explain the required field strength to account for the non-observation of GeV photons in blazar observations. Recently, it has been suggested that the reconnection time, rather than the Alfvén time, is the relevant timescale for the decay of the nonlinear evolution of magnetic fields in the early universe [94, 95, 96, 97]. In ref. [98], it was found that this makes the timescale approximately 50 times longer than the Alfvén time. If we account for this in our calculations, the length scale would decrease by a factor of $50^{2/3}$, and the magnetic field strength would increase by a factor of $50^{1/3}$ compared to the values mentioned in table 2. (If the decay is controlled by reconnection and the decay time still proportional to the square root of the conductivity, the exponents might be 4/7 and 2/7, respectively; see ref. [99].)

### v2

@@ 4 Gravitational waves
In this section, we discuss the production of GWs during axion-U(1) inflation. The interaction between the axion and the gauge field results in the generation of GWs, as the produced gauge field has a nonzero transverse traceless component in its energy-momentum tensor. Additionally, scalar field spatial fluctuations produced by the gauge field also contribute to the generation of GWs. GWs are defined as the transverse traceless part of the tensor perturbations ($h_{ij}^{\mathrm{TT}}$) of the metric. Including these tensor perturbations, the metric is given by
@@
The evolution of these tensor perturbations is governed by the Einstein equation. By linearizing the Einstein equation, we obtain
@@
Here, the ellipsis represents the irrelevant terms proportional to $g_{ij}$. The energy density of the produced GWs is given by
@@
The density fraction of GWs at the present epoch is given by
$\Omega_{\mathrm{GW}}h^{2}$ $=\frac{\rho_{\mathrm{GW}}}{\rho_{c}}\Big{|}_{0}h^{2}=\frac{\rho_{\mathrm{GW}}}{\rho_{r}}\Big{|}_{0}\Omega_{r}h^{2}=\frac{\rho_{\mathrm{GW}}|_{e}a_{e}^{4}}{\rho_{r}|_{0}a_{0}^{4}}\Omega_{r}h^{2}.$ (4.5)
In the above expression, we use $\rho_{r}|_{0}=\Omega_{r}\rho_{c}|_{0}$ and $\rho_{GW}|_{0}=\rho_{GW}|_{e}(a_{e}/a_{0})^{4}$. Here, $\Omega_{r}$ represents the radiation density fraction at the present epoch, $\rho_{GW}$ denotes the GW energy density at the epoch ‘e’ [^9] Here, ‘e’ refers to an epoch when the Universe is radiation-dominated and after this epoch, there is no further production of GWs. and $a_{e}$ and $a_{0}$ are the scale factors at epoch ‘e’ and the present epoch, respectively.. By considering the adiabatic evolution of the Universe, i.e., $g_{s}^{1/3}aT=\mbox{const}$, we obtain
$\Omega_{\mathrm{GW}}h^{2}$ $=\frac{\rho_{\mathrm{GW}}|_{e}g_{0,s}^{4/3}T_{0}^{4}}{\rho_{r}|_{0}g_{e,s}^{4/3}T_{e}^{4}}\Omega_{r}h^{2}.$ (4.6)
Here, $T_{0}$ and $T_{e}$ represent the temperature of the Universe at the present and at the epoch ‘e’, respectively. Furthermore, $g_{e,s}$ and $g_{0,s}$ denote the effective degrees of freedom in the entropy density at the epoch ‘e’ and the present epoch, respectively. Further, by using the radiation energy density at the present epoch, $\rho_{r}|_{0}=g_{0}\pi^{2}T_{0}^{4}/30$, and at the epoch ‘e’, $\rho_{r}|_{e}=g_{e}\pi^{2}T_{e}^{4}/30$, we obtain
$\Omega_{\mathrm{GW}}h^{2}$ $=\left(\frac{g_{0,s}}{g_{e,s}}\right)^{4/3}\frac{g_{e}}{g_{0}}\frac{\rho_{\mathrm{GW}}}{\rho_{r}}\Big{|}_{e}\Omega_{r}h^{2}\,$
$=0.4\times\left(\frac{g_{0,s}}{3.94}\right)^{4/3}\frac{g_{e}}{106.75}\left(\frac{g_{0}}{3.36}\right)^{-1}\left(\frac{g_{e,s}}{106.75}\right)^{-4/3}\frac{\rho_{\mathrm{GW}}}{\rho_{r}}\Big{|}_{e}\Omega_{r}h^{2}\,,$ (4.7)
where $g_{e}$ and $g_{0}$ represent the radiation degrees of freedom in energy density at the ‘e’ and present epochs, respectively. The above expression represent $\Omega_{\mathrm{GW}}h^{2}$ in terms of the ratio of the mean GW energy density and the total energy density at the epoch ‘e’, $\rho_{\mathrm{GW}}/\rho_{r}|_{e}$.
@@
Further in this section, we study the production of GWs in the axion-gauge field scenario, considering both weak and strong coupling cases. The characteristic current frequency of these produced GWs is in GHz range and is determined by the horizon scale at the end of inflation, see equation (5.3). In figure 3, we show the time evolution of the gauge field energy density normalized by the total energy density in the left panel and the density fraction of the GWs in the right panel. The curves correspond to different coupling strengths; $\alpha/f=30$ (dotted violet), 50 (dashed blue), 60 (dot-dashed blue), 75 (solid orange), and 90 (dot-dashed red). The vertical lines, in corresponding colors, represent the end of inflation for each case. The horizontal solid gray line indicates the bound on the GW energy density from the $\Delta N_{\mathrm{eff}}$ constraints from the CMB. The dashed gray curve represents the sensitivity of the future CMB-S4 mission. In the right panel of figure 3, we use the ratio of the GW energy density to the total energy density obtained from our simulations as $\rho_{\mathrm{GW}}/\rho_{r}|_{e}$ in equation (4) to determine the time-dependent value of $\Omega_{\text{GW}}$ shown on the $y$ axes. Our key findings regarding the constraints on the coupling strength between axion and gauge fields from the CMB bound on $\Delta N_{\mathrm{eff}}$ can be summarized as follows.
@@
  • As shown in figure 1, for $\alpha/f=75$ and $\alpha/f=90$, once the energy density of the gauge field becomes comparable to the kinetic energy of the axion field, the velocity of the axion decreases. This reduction in axion velocity leads to a slower growth rate of the gauge field compared to the pre-backreaction stage. Consequently, the production rate of GWs also slows down, as shown in figure 3. The amplitude of the generated GWs remains below the bound obtained using Planck data for these cases up until the end of inflation.
@@
  • Previous studies [38, 39] indicated that the coupling strength above $\alpha/f=70$ leads to an overproduction of GWs, violating the $\Delta N_{\mathrm{eff}}$ bound from the Planck data of CMB. This bound is for the case of chaotic potential, which is also considered in this work. However, we find that this bound is relaxed in the strong coupling case when gauge field backreaction becomes significant during inflation. The specific details of this bound depend on the duration of GW production after the end of inflation. As evident from figure 3, the produced GWs for $\alpha/f=75$ may violate the bound from the Planck data on the GW amplitude, but this occurs 2–3 $e$-folds after the end of inflation. In our analysis, we have not considered the interaction of the gauge field with other standard model particles. For large coupling strengths, this interaction becomes important and may lead to instantaneous reheating, as discussed in section 4.2 of ref. [49]. Considering these effects would lead to the decay of the gauge field into other standard model species, further affecting GW production. If the gauge field energy density decays significantly within 2–3 $e$-folds after inflation, the produced GW amplitude will be smaller and may well be within the bound obtained from the Planck data (shown by the solid horizontal gray line in figure 3).
    In addition, we observe that, for the cases $\alpha/f=75$ and $\alpha/f=90$, the growth rate of the gauge field energy density decreases at the time when the backreaction of the gauge field becomes significant. This is shown in the left panel of figure 3 as solid yellow and dash-dotted red curves. This allows more time for the production of charged particles compared to cases where backreaction becomes important only after inflation, such as $\alpha/f=50$ (dashed cyan curves) and $60$ (dash-dotted blue curves) shown in figure 3. From this discussion, we expect that the bound [38, 39] on $\alpha/f$ coming from Planck datasets will be relaxed. Indeed, assessing the status of the bound requires the inclusion of the charge current (which depends on charge particle content and high energy physics at energy scales $10^{14}$ GeV) which can further only suppress the production of the gauge fields and consequently GWs. A detailed study including the effects of charged particles is beyond the scope of this paper. We will report that in a future publication.
@@
    Figure 4: Time evolution of $\rho_{g}/\rho$ and $\Omega_{GW}h^{2}$ for $\alpha/f=75$ for run D (solid yellow curves) and for a run (dotted black curves) analogous to run D in which the contribution of the $\phi$ term in the total energy-momentum tensor, given in equation(4.3), is neglected in the calculation of the GW energy density.
@@ 5 Magnetogenesis
There is indirect evidence of magnetic fields in the intergalactic medium, inferred from the non-detection of secondary GeV photons from blazars [94, 95, 96, 97]. The presence of such intergalactic magnetic fields motivates the study of their generation in the early universe, potentially during inflation [98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108] or a first-order phase transition [109, 110, 111, 112, 113], for reviews see ref. [114, 115, 116]. In this section, we study the generation of magnetic fields within the axion-U(1) inflation scenario and calculate their present-day strength for different coupling strengths, $\alpha/f$.
@@ 5.1 Gauge-field production
First, we study the evolution of the electric and magnetic components of the gauge fields during inflation in our simulations. The root mean square value of the hypermagnetic field and its coherence length are given by [114, 115],
@@
Similar definitions are used for the electric part of the gauge fields. Furthermore, assuming the universe evolves in a radiation-dominated phase after epoch ‘e’ with no further magnetic field generation, we can calculate the magnetic field strength today using the following expression,
$B_{\mathrm{rms}}\big{|}_{0}$ $=B_{rms}|_{e}\left(\frac{a_{e}}{a_{0}}\right)^{2}=B_{rms}|_{e}\left(\frac{g_{0,s}^{2/3}T_{0}^{2}}{g_{e,s}^{2/3}T_{e}^{2}}\right)=\sqrt{\frac{{\int d\log{k}\cdot P_{B}}}{\rho_{e}}}\sqrt{\frac{\pi^{2}g_{e}}{30}}\left(\frac{g_{0,s}}{g_{e,s}}\right)^{2/3}T_{0}^{2}.$
$=1.9\,{\rm\mu G}\,\sqrt{\frac{{\int d\log{k}\cdot P_{B}}}{\rho_{e}}}\left(\frac{g_{e}}{106.75}\right)\left(\frac{g_{0,s}}{3.94}\right)^{2/3}\left(\frac{106.75}{g_{e,s}}\right)^{2/3}\left(\frac{T_{0}}{2.73{\rm K}}\right)^{2}$ (5.2)
In the above expression, we assume the adiabatic evolution of the Universe, given by $g_{s}^{1/3}aT=const$, and use $\rho_{e}=g_{e}\pi^{2}T_{e}^{4}/3$. Here, $\rho_{e}$ represents the total energy density of the universe at epoch ‘e’, and $T_{0}$ denotes the CMB temperature.
In figure 5, we demonstrate the evolution of the electric component with solid lines and the magnetic component with dashed lines for coupling strengths $\alpha/f=30$ (green curve), $50$ (red curve), $60$ (orange curve), $75$ (blue curve), and $90$ (yellow curve). The values of $E_{\mathrm{rms}}$ and $B_{\mathrm{rms}}$ on the $y$-axis in the left panel are determined using the ratio $\sqrt{\int d\log{k}\cdot P_{B}/\rho_{e}}$ obtained from the simulation in equation (5.1). In the right panel, we illustrate the evolution of the coherence length for each coupling strength. The coherence length scale is normalized by the length scale corresponding to the comoving Hubble rate, $\mathcal{H}$, as a function of time. The length scale corresponding to the Hubble horizon at the end of inflation is given by
@@
$\frac{a_{e}}{a_{0}}=\left(\frac{g_{0,s}}{g_{e,s}}\right)^{1/3}\frac{T_{0}}{T_{e}}=2.66\times 10^{-29}\left(\frac{g_{0,s}}{3.94}\frac{106.75}{g_{e,s}}\right)^{1/3}\left(\frac{106.75}{g_{e}}\right)^{1/4}\frac{T_{0}}{2.73\text{K}}\left(\frac{10^{-6}m_{\rm pl}}{H}\right)^{1/2},$ (5.4)
where $g_{e,s}$ and $g_{0,s}$ denote the effective degrees of freedom in the entropy density at the end of inflation and the present epoch, respectively. We estimate the reheating temperature, $T_{r}$, by assuming instantaneous reheating using $3H^{2}m_{\text{pl}}^{2}=8\pi(\pi^{2}/30)g_{e}T_{e}^{4}$.
Figure 5: The evolution of the electric (solid lines) and magnetic (dashed lines) components of the gauge field (left panel) and their coherence lengths (right panel) is shown for different values of $\alpha/f$. The coherence lengths are normalized by the length scale corresponding to the comoving Hubble rate ($\mathcal{H}=a^{\prime}/a$) as a function of time.
@@ 5.2 Evolution of gauge fields post inflation
After the end of inflation, the universe is dominated by the gauge field in the strong coupling regime of axion-U(1) system. Due to the dominance of the gauge field in the total energy budget, the effective equation of state resembles that of radiation dominance. We assume that reheating in this model occurs through the interaction of the gauge field with other Standard Model (SM) fields, as described in [122, 49]. Therefore, we can consider the universe as being radiation-dominated just after the end of inflation.
As discussed earlier, the gauge fields have a blue spectrum at the end of inflation and peaks at a scale smaller than the Hubble horizon size; see figure 2. We assume that the magnetic part of the gauge fields remain unaffected  [^10] Thus, we effectively assume that only the electric part of the gauge fields energy gets transferred to the SM particles. This is a simplified scenario. In particular one can expect that an important role should be played by the charge currents. Moreover, depending on the reheating scenario magnetic energy may also be reduced. However, the exact mechanism of reheating and energy transfer lies outside of the scope of this work. from the end of inflation until the epoch when the universe’s conductivity becomes much larger than the Hubble parameter. Once this epoch is reached, the electric component of the gauge field shortened out within a Hubble time scale, and the magnetic field becomes frozen into the rest of the plasma. After this epoch, the magnetic field interacts with the plasma via the Lorentz force and transfers some of its energy to the plasma. As a result, the magnetic field strength decreases. This nonlinear processing of the magnetic field begins when the Alfvén time becomes smaller than the Hubble time (see section 6 of ref. [115] for details), which implies that the corresponding scale factor is $a_{\textrm{nl}}\equiv a_{H}(\mathcal{H}~{}L_{c}/V_{A})$. Here, $a_{H}$ corresponds to the scale factor at the horizon entry for a particular scale, and $V_{A}=\sqrt{B^{2}/(\rho+p)}$ is the Alfvén velocity.
@@
A $35$ $6.2\times 10^{-17}$ $2.0\times 10^{-3}$ $1.8\times 10^{-13}$ $1.1\times 10^{-15}$ $1.3\times 10^{-2}$ $7.0\times 10^{-14}$
B $50$ $4.5\times 10^{-16}$ $1.2\times 10^{-2}$ $7.3\times 10^{-14}$ $1.3\times 10^{-14}$ $1.5\times 10^{-1}$ $2.1\times 10^{-14}$
C $60$ $1.7\times 10^{-15}$ $4.1\times 10^{-2}$ $4.0\times 10^{-14}$ $2.7\times 10^{-14}$ $1.5\times 10^{-1}$ $2.1\times 10^{-14}$
D $75$ $1.5\times 10^{-14}$ $2.5\times 10^{-1}$ $1.6\times 10^{-14}$ $2.6\times 10^{-14}$ $2.4\times 10^{-1}$ $1.7\times 10^{-14}$
E $90$ $1.9\times 10^{-14}$ $2.5\times 10^{-1}$ $1.6\times 10^{-14}$
Table 2: $B_{\mathrm{rms}}|_{0}$ and $L_{c}^{\mathrm{NL}}|_{0}$ represent the present-day magnetic field strength and its coherence length obtained for each run, respectively. These $B_{\mathrm{rms}}|_{0}$ and $L_{c}^{\mathrm{NL}}|_{0}$ in columns 3 and 4 are calculated from the spectrum at the end of inflation. For the sake of comparison, in columns 6 and 7 we provide $B_{\mathrm{rms}}^{1}|_{0}$ and $L_{c}^{1,\mathrm{NL}}|_{0}$ calculated from the spectrum obtained one $e$-fold after the end of inflation, provided we have these data. $B_{\rm bound}$ denotes the lower bound on the magnetic field strength based on blazar observations, where we assume $B_{\rm bound}=1.8\times 10^{-17}{\rm G}\left(L/0.2\text{Mpc}\right)^{1/2}$ [123].
In the nonlinear processing regime, the magnetic field strength decays as $(a_{m}/a_{nl})^{-1/3}$, and the coherence length (which corresponds to the peak scale in our case) increases as $(a_{m}/a_{nl})^{2/3}$ for the fully helical case. [^11] As demonstrated in appendix D for $\alpha/f=75$ and $\alpha/f=90$, the relative magnetic helicity near the end of inflation depends on the wave number. Notably, the field remains almost fully helical around the peak wave number of the power spectrum of the magnetic component. Hence, a similar late-time evolution is expected as in the fully helical case. However, a detailed investigation would require MHD simulations in the radiation-dominated epoch for treating such a magnetic field energy density spectrum. By using $T=1{\rm eV}$ (the temperature of the Universe at the matter-radiation equality) in equation (5.4), we obtain,
$\frac{a_{nl}}{a_{m}}=1.13\times 10^{-25}\left(\frac{g_{m,s}}{3.94}\frac{106.75}{g_{nl,s}}\right)^{1/3}\left(\frac{106.75}{g_{nl}}\right)^{1/4}\frac{T_{m}}{1{\rm eV}}\left(\frac{10^{-6}m_{\rm pl}}{H}\right)^{1/2}r.$ (5.5)
Here $r=\max(1,\mathcal{H}~{}L_{c}/V_{A})$ and $T_{m}$ denotes the temperature of the Universe at the matter-radiation equality. We show the time evolution of $\mathcal{H}L_{c}/V_{A}$ in appendix E for the runs given in table 1. After incorporating the nonlinear processing of the fields, the magnetic field strength and its coherence length at the present epoch is given by
$B_{\mathrm{rms}}\big{|}_{0}$ $=9.2\times 10^{-15}\,{\rm G}\,\sqrt{\frac{{\int d\log{k}\cdot P_{B}}}{\rho_{e}}}\left(\frac{10^{-6}m_{\rm pl}}{H}\right)^{1/6}r^{1/3}$ (5.6)
@@
$L_{c}^{\mathrm{NL}}=0.8~{}{\rm pc}~{}(\mathcal{H}L_{c})\left(\frac{10^{-6}m_{\rm pl}}{H}\right)^{1/6}\,r^{-2/3}.$ (5.7)
@@
After using the above expressions, we obtained the magnetic field strength and its coherence length corresponding to each coupling strength and the values are given in table 2. For the values given in columns 3 and 4 of table 2, we use $({\int d\log{k}\cdot P_{B}}/\rho)^{1/2}$ and $\mathcal{H}L_{c}$ at the end of inflation in equation (5.6) and equation (5.7), respectively. However, for the values given in columns 6 and 7, we use the value of $({\int d\log{k}\cdot P_{B}}/\rho)^{1/2}$ and $\mathcal{H}L_{c}$ one $e$-fold after inflation for the sake of comparison with the numbers in columns 2 and 3. For run E, we do not provide values at one $e$-fold after inflation, as this would require a simulation with a resolution greater than $1024^{3}$. In table 1, $B_{\rm bound}$ denotes the lower bound of magnetic field strength obtained from the blazars observation. Here we assume $B_{\rm bound}=1.8\times 10^{-17}{\rm G}\,(L_{C}^{\mathrm{NL}}/0.2{\rm Mpc})^{1/2}$ [123]. From the values given in the table 2, we conclude that the magnetic field strength obtained in the axion-U(1) inflation model for coupling strengths $\alpha/f\geq 60$ can explain the required field strength to account for the non-observation of GeV photons in blazar observations. Recently, it has been suggested that the reconnection time, rather than the Alfvén time, is the relevant timescale for the decay of the nonlinear evolution of magnetic fields in the early universe [124, 125, 126, 127]. In ref. [128], it was found that this makes the timescale approximately 50 times longer than the Alfvén time. If we account for this in our calculations, the length scale would decrease by a factor of $50^{2/3}$, and the magnetic field strength would increase by a factor of $50^{1/3}$ compared to the values mentioned in table 2. (If the decay is controlled by reconnection and the decay time still proportional to the square root of the conductivity, the exponents might be 4/7 and 2/7, respectively; see ref. [129].) In figure 6, we illustrate the obtained magnetic field strength and its coherence length for different values of $\alpha/f$, along with the lower and upper limits on the magnetic field strength from various observations. The purple, cyan, green, black, and red points correspond to $\alpha/f=35,50,60,75$, and 90, respectively, in the case where the Alfvén time is taken as the decay time scale. The stars represent the corresponding values when the reconnection time is considered as the decay time scale.
Figure 6: Obtained magnetic field strength versus its coherence length for different values of $\alpha/f$. The purple, cyan, green, black, and red correspond to $\alpha/f=35,50,60,75,90$, respectively. Bullet points, correspond to the case in which Alfvén time is considered as the decay timescale for the nonlinear evolution of the magnetic field after its generation. However, if one considers the reconnection time instead, the values change, and these are represented by stars in the corresponding colors. The black-shaded region denotes the lower bound on the inter-galactic magnetic field strength obtained from GeV observations of blazars [123], while the red region represents the upper bound from Big Bang Nucleosynthesis (BBN) [130]. The blue region indicates the upper bound obtained from Planck collaboration [131] and blue dotted line represents the improved upper bound considering inflationary magnetic curvature mode in non-Gaussianity trispectrum [132].
In our study, we have not considered the production of fermions resulting from interactions between gauge fields and fermions, e.g., via the Schwinger effect. This has been discussed in refs. [85, 53] for the case of a homogeneous axion field. Upon incorporating the effects of charged currents generated by the Schwinger effect, the strength of the produced magnetic fields is expected to decrease by a factor of less than about 10, as can be seen from figures 6 and 7 of ref. [53]. However, the exact reduction factor depends on the coupling strength. The impact of this effect in the case of an inhomogeneous axion field is not known, and we plan to address this in future work. In summary, we note that the gauge field strengths presented in table 2 could be reduced if the effect of the produced charged currents were included.